OpenStudy (anonymous):

184x + 4 y -64 =0 written as a function of x would it be f (x) = 16 + 46 x???

OpenStudy (anonymous):

184x+4y-64=0 4y=64-184x y=16-46x f(x)=16-46x

OpenStudy (anonymous):

thx for writing it down it helps

OpenStudy (anonymous):

16+46x?? =>184x+4y-64=0 4y=64-184x y=16-46x

OpenStudy (anonymous):

yea thx to u too Rohangrr